A composite wall is formed of a 2.5 cm copper plate (k=385 W/m-K). a 3.2 cm layer of

asbestos (k=0.2 W/m-K), and a 5 cm layer of fibreglass (k=0.045 W/m-K). The wall is subjected to an overall temperature drop of 560
1
Expert's answer
2021-08-20T01:36:45-0400

thickness of copper plate = 2.5cm

kc = 385 W/m.K

thickness of absestos = 3.2cm

ka = 0.2 W/m.K

thickness of fireglass = 5cm

kf = 0.045 W/m.K

∆T = 560°

rate of heat loss (Q) =


RthC = "\dfrac{L_c}{k_cA} = \dfrac{2.5×10^{-2}}{385×1}"


"= 6.5\u00d7 10^{-5}\u00b0C\/W"


RthA = "\dfrac{L_A}{k_AA} = \dfrac{3.2×10^{-2}}{0.2×1}"


"=0.16\u00b0C\/W"


RthF = "\dfrac{L_F}{k_FA} = \dfrac{5×10^{-2}}{0.045×1}=1.11°C/W"


"Q = \dfrac{∆T}{R_{thC}+R_{thA}+R_{thF}}="


"\dfrac{560}{6.5×10^{-5}+0.16+1.11}= 440.92\ W"